OpenStudy (baseballer2014):

Calculus Sequence Help!! What is the general term for the sequence 1,2,1,2,1...

ganeshie8 (ganeshie8):

\[\frac{3}{2} \color{Red}{-} \frac{1}{2},~~\frac{3}{2} \color{red}{+} \frac{1}{2},~~\frac{3}{2} \color{Red}{-} \frac{1}{2},~~\frac{3}{2} \color{red}{+} \frac{1}{2},~\cdots \] can you write the generall term of this sequence ?